# Quantum gravity inspired nonlocal quantum dynamics preserving the classical limit

### v1

@@ II.1 Non-relativistic limit of nonlocal effective field theories
The nonrelativistic complex field $\psi$ satisfies a nonlocal generalisation of the Schrödinger equation. In order to interpret $\psi$ as the wavefunction of a quantum mechanical system one should construct from the field operator $\psi$ a wavefunction for a generic one particle state and show that it satisfies the same Schrödinger equation as the field. While this is indeed the case at the first-order of a perturbative expansion of $f(S)$ belen19 , the generalization of this result to the full theory would require the derivation of the nonlocal Hamiltonian, a result that is technically challenging. Thus, from here on we will proceed with the caveat that our approach is phenomenological, in the sense that a rigorous identification of $\psi$ with the wavefunction of a quantum mechanical system is yet to be established.
@@ II.2 Non-local Schrödinger evolution in a harmonic potential
where $m$ is the oscillator mass, $\omega$ its resonant frequency and $\epsilon=2ml_{k}^{2}/\hbar^{2}$. Using dimensionless coordinates $\hat{t}=\omega t$ and $\hat{x}=x/x_{0}$, Eq. 4 takes the form
@@
where $\hat{\mathcal{S}}=i\partial_{t}+\frac{1}{2}\partial_{\hat{x}\hat{x}}^{2}$ and $\hat{\epsilon}=\epsilon\hbar\omega$. Eq. (5) reduces to the standard Schrödinger equation in the local limit $\hat{\epsilon}=0$. Note that $\sqrt{\hat{\epsilon}}$ corresponds to the ratio between $l_{k}$ and the width of the oscillator’s ground-state wavefunction $x_{0}=\sqrt{\hbar/2m\omega}$ or, equivalently, $\hat{\epsilon}$ gives the ratio between the quantum ground state energy and the nonlocality energy scale. Such dependence suggests that massive quantum systems or, more precisely, systems with the smallest zero-point fluctuations could be the ideal setting for detecting these deviations. For a nonlocality length ranging between $l_{k}\sim 10^{-19}$m (higher values have been already excluded biswas ) and the Planck scale $l_{k}\sim 10^{-35}$m, a quantum oscillator with $x_{0}=10^{-18}$m (see e.g. Ref. bild ) would experience deviations from the ordinary Schrödinger evolution ranging from $\hat{\epsilon}\sim 10^{-2}$ to $\hat{\epsilon}\sim 10^{-34}$. From now on, for simplicity, we shall omit hats in all quantities.
@@
where $k$ is the wavenumber, $\tilde{\phi}(k)$ is the Fourier transform of $\phi(x)$ and $V_{nl}(E,k)=E+(\frac{1}{2}k^{2}-E)e^{-\epsilon(\frac{k^{2}}{2}-E)}$. Hence, the nonlocal quantum evolution in a harmonic potential in real space is mapped into a standard, local quantum dynamics in an energy-dependent potential. The latter can be seen as encoding the residual backreaction of the quantum field on the background space. Interestingly, in the case of a quantum harmonic oscillator, the above correspondence holds for any analytic function $f$.
@@ III.2 Energy spectrum and quantum probability densities
Figure 1: a) Energy eigenvalues $E_{n}$ obtained from numerical integration of Eq. 7 for $\epsilon=10^{-7}$ (black stars), $\epsilon=10^{-4}$ (red diamonds), $\epsilon=3\times 10^{-4}$ (blue circles), $\epsilon=\times 10^{-3}$ (magenta squares), $\epsilon=3\times 10^{-3}$ (green triangles) and the corresponding fitting curves (solid lines) using Eq. (14) with $a$ as free parameter. b) Best-fit values of the parameter $a$ for different $\epsilon$. c) Collapse of the numerical data shown in a) onto the scaling function $E^{\prime}\ln(1+n^{\prime})$ (solid black curve), where $E^{\prime}=\epsilon a(E_{n}-1/2)$ and $n^{\prime}=\epsilon an$. Additional data corresponding to other $\epsilon$ (see panel b)) are depicted in grey.
@@ IV Nonlocal Helmholtz free energy
$\sum_{n=0}^{\infty}G(n)=\sum_{n=0}^{\infty}(1+a\epsilon n)^{-\frac{\beta}{\epsilon a}}=\frac{1}{\beta}$ (19)
Let us thus look at the convergence of the series on the left hand side of Eq. (19). The integral test for convergence applied to the infinite series gives the lower and upper bounds
[…]
from which we derive the lower bound:
[…]
Since $G(y=0)=1$ we finally obtain
$\frac{1}{\beta-\epsilon a}\leq\sum_{n=0}^{\infty}G(n)\leq 1+\frac{1}{\beta-\epsilon a}.$ (23)
In the regime in which $1\ll 1/\beta\ll 1/\epsilon$, i.e. for thermal energies much larger than the zero-point energy, but also much smaller that the nonlocality energy scale, Eq. (23) implies $\sum_{n=0}^{\infty}G(n)=1/\beta$. Thus, we see that we retrieve the classical limit for the Helmholtz free energy of a local quantum harmonic oscillator.

### v2

@@ II.1 Non-relativistic limit of nonlocal effective field theories
The nonrelativistic complex field $\psi$ satisfies a nonlocal generalisation of the Schrödinger equation. In order to interpret $\psi$ as the wavefunction of a quantum mechanical system one should construct from the field operator $\psi$ a wavefunction for a generic one particle state and show that it satisfies the same Schrödinger equation as the field. While this is indeed the case at the first-order of a perturbative expansion of $f(\mathcal{S})$ belen19 , the generalization of this result to the full theory would require the derivation of the nonlocal Hamiltonian, a result that is technically challenging. Thus, from here on we will proceed with the caveat that our approach is phenomenological, in the sense that a rigorous identification of $\psi$ with the wavefunction of a quantum mechanical system is yet to be established.
@@ II.2 Non-local Schrödinger evolution in a harmonic potential
where $m$ is the oscillator mass, $\omega$ its resonant frequency and $\epsilon=2ml_{k}^{2}/\hbar^{2}$. Notice that $\epsilon$ has dimensions of the inverse of an energy. By introducing the nonlocality energy scale as $E_{k}=\hbar c/l_{k}$, from the above definition of $\epsilon$ we see that $\epsilon E_{k}=2mc^{2}/E_{k}$. The dimensionless quantity $\epsilon E_{k}$ gives (up to a factor 2) the ratio between the mass energy of the harmonic oscillator and the nonlocality energy scale. Thus $1/\epsilon$ is an energy scale directly related to the nonlocality energy scale when probed by an oscillator of mass $m$. For brevity, we will refer to $1/\epsilon$ simply as the nonlocality energy scale from now on.
In order to cast Eq. (4) in a dimensionless form amenable to theoretical and numerical analysis, we rescale the physical time as $\hat{t}=\omega t$, where $\omega$ is the angular frequency of the mechanical oscillator. The spatial coordinate $x$ is rescaled to $\hat{x}=x/x_{0}$, where $x_{0}=\sqrt{\hbar/m\omega}$. Using the dimensionless coordinates $\hat{t}=\omega t$ and $\hat{x}=x/x_{0}$, Eq. (4) takes the form
@@
where $\hat{\mathcal{S}}=i\partial_{\hat{t}}+\frac{1}{2}\partial_{\hat{x}\hat{x}}^{2}$ and $\hat{\epsilon}=\epsilon\hbar\omega$. Eq. (5) reduces to the standard Schrödinger equation in the local limit $\hat{\epsilon}=0$. Note that $\sqrt{\hat{\epsilon}}$ corresponds to the ratio between $l_{k}$ and the amplitude of zero-point fluctuations $x_{ZPF}=x_{0}/\sqrt{2}=\sqrt{\hbar/2m\omega}$ or, equivalently, $\hat{\epsilon}$ gives the dimensionless ratio between an energy quantum of the harmonic oscillator $\hbar\omega$ and the nonlocality energy scale $1/\epsilon$. Such dependence (as well as the above physical interpretation of $1/\epsilon$) suggests that massive quantum systems or, more precisely, systems with the smallest zero-point fluctuations could be the ideal setting for detecting these deviations. For a nonlocality length ranging between $l_{k}\sim 10^{-19}$m (higher values have been already excluded biswas ) and the Planck scale $l_{k}\sim 10^{-35}$m, a quantum oscillator with $x_{0}=10^{-18}$m (see e.g. Ref. bild ) would experience deviations from the ordinary Schrödinger evolution ranging from $\hat{\epsilon}\sim 10^{-2}$ to $\hat{\epsilon}\sim 10^{-34}$. From now on, for simplicity, we shall omit hats in all quantities.
@@
where $k$ is the dimensionless wavenumber ($k\rightarrow k\,x_{0}$), $\tilde{\phi}(k)$ is the Fourier transform of $\phi(x)$ and $V_{nl}(E,k)=E+(\frac{1}{2}k^{2}-E)e^{-\epsilon(\frac{k^{2}}{2}-E)}$. Hence, the nonlocal quantum evolution in a harmonic potential in real space is mapped into a standard, local quantum dynamics in an energy-dependent potential. The latter can be seen as encoding the residual backreaction of the quantum field on the background space. Interestingly, in the case of a quantum harmonic oscillator, the above correspondence holds for any analytic function $f$.
@@ III.2 Energy spectrum and quantum probability densities
Figure 1: a) Energy eigenvalues $E_{n}$ obtained from numerical integration of Eq. 7 for $\epsilon=10^{-7}$ (black stars), $\epsilon=10^{-4}$ (red diamonds), $\epsilon=3\times 10^{-4}$ (blue circles), $\epsilon=10^{-3}$ (magenta squares), $\epsilon=3\times 10^{-3}$ (green triangles) and the corresponding fitting curves (solid lines) using Eq. (14) with $a$ as free parameter. b) Best-fit values of the parameter $a$ for different $\epsilon$. c) Collapse of the numerical data shown in a) onto the scaling function $E^{\prime}\ln(1+n^{\prime})$ (solid black curve), where $E^{\prime}=\epsilon a(E_{n}-1/2)$ and $n^{\prime}=\epsilon an$. Additional data corresponding to other $\epsilon$ (see panel b)) are depicted in grey.
@@ IV Nonlocal Helmholtz free energy
$\sum_{n=0}^{\infty}(1+a\epsilon n)^{-\frac{\beta}{\epsilon a}}=\frac{1}{\beta}$ (19)
The integral test for convergence applied to the series on the left hand side of Eq. (19) gives the lower and upper bounds (see Appendix)
$\frac{1}{\beta-\epsilon a}\leq\sum_{n=0}^{\infty}(1+a\epsilon n)^{-\frac{\beta}{\epsilon a}}\leq 1+\frac{1}{\beta-\epsilon a}.$ (20)
In the regime in which $1\ll 1/\beta\ll 1/\epsilon$, i.e. for thermal energies much larger than the zero-point energy, but also much smaller that the nonlocality energy scale, Eq. (20) implies $\sum_{n=0}^{\infty}(1+a\epsilon n)^{-\frac{\beta}{\epsilon a}}=1/\beta$. Thus, we see that we retrieve the classical limit for the Helmholtz free energy of a local quantum harmonic oscillator.
@@ VI Conclusions
Acknowledgments
AB acknowledges support from the Horizon Europe EIC Pathfinder project QuCoM (Grant Agreement No. 10104697).
Appendix
Let us thus look at the convergence of the series
$\sum_{n=0}^{\infty}G(n)=\sum_{n=0}^{\infty}(1+a\epsilon n)^{-\frac{\beta}{\epsilon a}}=\frac{1}{\beta}$
The integral test for convergence applied to the infinite series gives the lower and upper bounds
[…]
from which we derive the lower bound
[…]
Since $G(y=0)=1$ we finally obtain
$\frac{1}{\beta-\epsilon a}\leq\sum_{n=0}^{\infty}G(n)\leq 1+\frac{1}{\beta-\epsilon a}.$